# Protect Measurement-Induced Phase Transition from Noise

### v1

@@ Abstract
$\sum_{m}\Lambda_{A}^{Q}$ $=$ $\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{A})=\text{Tr}_{A\cup E}\left(\text{SWAP}^{\prime}\rho^{\otimes Q}\text{SWAP}^{\prime\dagger}\right.$ (5)
$\left.\mathbb{C}^{A}\right)=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{E})\equiv\sum_{m}\Lambda^{Q}_{E}.$
@@
Discussions. Similar to QEC, where more physical qubits are required to encode a smaller number of logical qubits, our scheme necessitates the preparation of additional ancilla qubits as the circuit depth increases. Specifically, $O(qNT)$ ancilla qubits are needed, where $N$ is the number of system qubits, $p$ is the noise rate, and $T$ is the circuit depth. Assuming $q_{e}\sim 1\%,L=20$ and $T=10L$, around $800$ ancilla qubits are needed to observe the phase transition. It is important to highlight that each ancilla qubit is utilized only once and can be isolated after use, thereby mitigating the risk of noise contamination.
An immediate application of our scheme requires prior knowledge of the probability that a specific type of decoherence noise will occur in the experimental setting. Our scheme can be directly adapted to determine the noise rate as follows: apply a finite amount of measurement such that the state enters the area-law phase, then sweep through different QE operation rates $q_{e}$. In a finite-size system, $\overline{S(M|A)}$ will be greater than zero if $q_{e}>q_{n}$ and less than zero otherwise.

### v2

@@ Abstract
$\sum_{m}\Lambda_{A}^{Q}$ $=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{A})=\text{Tr}_{A\cup E}\left(\text{SWAP}^{\prime}\rho^{\otimes Q}\right.$ (5)
$\left.\text{SWAP}^{\prime\dagger}\mathbb{C}^{A}\right)=\text{Tr}_{A\cup E}(\rho^{\otimes Q}\mathbb{C}^{E})\equiv\sum_{m}\Lambda^{Q}_{E}.$
@@
Discussions. Similar to QEC, where more physical qubits are required to encode a smaller number of logical qubits, our scheme necessitates the preparation of additional ancilla qubits as the circuit depth increases. Specifically, $O(qNT)$ ancilla qubits are needed, where $N$ is the number of system qubits, $p$ is the noise rate, and $T$ is the circuit depth. Assuming $q_{e}\sim 1\%,L=20$ and $T=10L$, around $800$ ancilla qubits are needed to observe the phase transition. It is imperative to underscore that we ignore potential noise in the ancilla qubits. This is justified, as each ancilla qubit is utilized only once and can be effectively isolated after use, thereby mitigating the risk of noise propagation and contamination.
An immediate application of our scheme requires prior knowledge of the probability that a specific type of decoherence noise will occur in the experimental setting. Our scheme can be directly adapted to determine the noise rate as follows: apply a finite amount of measurement such that the state enters the area-law phase, then sweep through different QE operation rates $q_{e}$. In the thermodynamic limit, curves corresponding to different $p$ would intersect at the point $q_{n}=q_{e}$. A more detailed explanation and finite size numerical results are provided in [46].